# InfAlign: Inference-aware language model alignment

### v2

@@ A.2 Proof of Theorem 1
The proof follows from the following identities:
$W^{T}_{r}(\pi_{1}\succ\pi_{2}\mid{\bm{x}})$ $=\mathbb{E}_{{\bm{z}}\sim T(\pi_{1})(\cdot\mid{\bm{x}}),{\bm{y}}\sim T(\pi_{2})(\cdot\mid{\bm{x}})}\left\{w_{r}({\bm{z}},{\bm{y}}\mid{\bm{x}})\right\}$
$=\mathbb{E}_{{\bm{z}}\sim T(\pi_{1})(\cdot\mid{\bm{x}})}\left\{\mathbb{E}_{{\bm{y}}\sim T(\pi_{2})(\cdot\mid{\bm{x}})}\left\{w_{r}({\bm{z}},{\bm{y}}\mid{\bm{x}})\right\}\right\}$
$=\mathbb{E}_{{\bm{z}}\sim T(\pi_{1})(\cdot\mid{\bm{x}})}\left\{\mathcal{C}_{r,T(\pi_{2})}({\bm{x}},{\bm{z}})\right\}$
$=\sum_{{\bm{z}}}\mathcal{C}_{r,T(\pi_{2})}({\bm{x}},{\bm{z}})T(\pi_{1})({\bm{z}}\mid{\bm{x}}).$
@@
$\mathcal{L}(\pi(\cdot\mid{\bm{x}}),\alpha)=W^{T}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})-\beta D_{\text{KL}}(\pi(\cdot\mid{\bm{x}})\|{\pi_{\rm ref}}(\cdot|{\bm{x}}))+\alpha\left(\sum_{{\bm{y}}}\pi({\bm{y}}\mid{\bm{x}})-1\right).$
@@
$\mathbf{0}=\frac{\partial\mathcal{L}(\pi(\cdot\mid{\bm{x}}),\alpha)}{\partial\pi({\bm{y}}\mid{\bm{x}})}=\frac{\partial W^{T}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})}{\partial\pi({\bm{y}}\mid{\bm{x}})}-\beta\left(\log\frac{\pi({\bm{y}}\mid{\bm{x}})}{{\pi_{\rm ref}}({\bm{y}}\mid{\bm{x}})}+1\right)+\alpha.$
@@
$\mathcal{R}({\bm{x}},{\bm{y}})=\frac{\partial W^{T}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})}{\partial\pi({\bm{y}}\mid{\bm{x}})},$
@@
It remains to prove Eq. 9. Plugging in $\pi_{1}=\pi$, $\pi_{2}={\pi_{\rm ref}}$ to Lemma 7, and taking partial derivative with respect to $\pi({\bm{y}}\mid{\bm{x}})$ on the right hand side completes the proof. ∎
A.3 Proof of Lemma 2
@@ A.5 Proof of Lemma 4
$\text{Pr}_{{\bm{y}}\sim{\pi_{\rm ref}}}\left(\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\leq u\right)$ $=\text{Pr}_{{\bm{y}}\sim{\pi_{\rm ref}}}{r({\bm{x}},{\bm{y}})\leq r({\bm{x}},{\bm{y}}_{u,{\bm{x}}})}$
@@
A.6 Proof of Theorem 2
@@
If $T$ is a calibrated inference-time procedure with mapping function $g_{T}$, for any continuous language model $\pi$, $\beta>0$ and reward transformation function $\Phi$, we have that
$W^{T}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})=\frac{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{T}(F_{\Phi,\beta}(u))\int_{0}^{u}g_{T}(u^{\prime}){\rm d}u^{\prime}{\rm d}u}{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{T}(F_{\Phi,\beta}(u)){\rm d}u\int_{0}^{1}g_{T}(u){\rm d}u}$
@@
$\mathcal{C}_{r,{\pi_{\rm ref}}}\circ\pi^{*}_{\mathcal{R}_{\Phi},\beta}(u)=\frac{e^{\Phi(u)/\beta}}{\int_{0}^{1}e^{\Phi(u^{\prime})/\beta}{\rm d}u^{\prime}}.$
@@
After the inference-time procedure $T$ is applied, we have that inference-time base policy satisfies
$\mathcal{C}_{r,{\pi_{\rm ref}}}\circ T({\pi_{\rm ref}})(u)=\frac{g_{T}(u)}{\int_{0}^{1}g_{T}(u^{\prime}){\rm d}u^{\prime}}.$
@@
$\mathcal{C}_{r,{\pi_{\rm ref}}}\circ T(\pi^{*}_{\mathcal{R}_{\Phi},\beta})(u)=\frac{\exp\left(\Phi(u)/\beta\right)g_{T}(F_{\Phi,\beta}(u))}{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{T}(F_{\Phi,\beta}(u)){\rm d}u}.$
Thus, the inference-time win rate satisfies [^4] Below we add an additional $\frac{1}{2}\mathbbm{1}\left\{r({\bm{x}},{\bm{y}})=r({\bm{x}},{\bm{z}})\right\}$ to the win r.v for simplicity, which won’t affect the result for continuous LMs.
$W^{T}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})$ $=\mathbb{E}_{{\bm{y}}\sim T(\pi^{*}_{\mathcal{R}_{\Phi},\beta})(\cdot\mid{\bm{x}}),{\bm{z}}\sim T({\pi_{\rm ref}})(\cdot\mid{\bm{x}})}\left\{\mathbbm{1}\left\{r({\bm{x}},{\bm{y}})\geq r({\bm{x}},{\bm{z}})\right\}\right\}$
$=\mathbb{E}_{{\bm{y}}\sim T(\pi^{*}_{\mathcal{R}_{\Phi},\beta})(\cdot\mid{\bm{x}}),{\bm{z}}\sim T({\pi_{\rm ref}})(\cdot\mid{\bm{x}})}\left\{\mathbbm{1}\left\{\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\geq\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{z}})\right\}\right\}$
$=\text{Pr}_{u\sim\mathcal{C}_{r,{\pi_{\rm ref}}}\circ T(\pi^{*}_{\mathcal{R}_{\Phi},\beta}),u^{\prime}\sim\mathcal{C}_{r,{\pi_{\rm ref}}}\circ T({\pi_{\rm ref}})}\left(u^{\prime}\leq u\right)$
$=\frac{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{T}(F_{\Phi,\beta}(u))\int_{0}^{u}g_{T}(u^{\prime}){\rm d}u^{\prime}{\rm d}u}{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{T}(F_{\Phi,\beta}(u)){\rm d}u\int_{0}^{1}g_{T}(u){\rm d}u},$
completing the proof. ∎
A.7 Proof of Theorem 3
@@ A.8 Proof of Corollary 2
Let ${\bm{\mathcal{Y}}}=[0,1]$. We assume the LMs and reward models are context-independent. We use $u\in[0,1]$ to denote ${\bm{y}}$ and set the reward model to be $r(u)=u$. The base policy is a simple uniform distribution over $[0,1]$, ${\pi_{\rm ref}}=\text{Unif}([0,1])$. Let $F_{\pi}(u)$ be the CDF of $\pi$, then we have that the BoN win rate is
@@
$=-N\int_{0}^{1}v^{N-1}F_{\pi}(v)^{N-1}\frac{\partial F_{\pi}(v)}{\partial\pi(u)}{\rm d}v$
@@
Appendix B The role of KL divergence in model alignment
@@
$\pi_{\theta}(y|x)=e^{\theta^{T}g(x,y)-A(\theta;x)},$ (18)
where $g(x,y)$ is a fixed encoding of $(x,y)$, and $A(\theta;x)$ is the partition function normalizing the distribution.
@@
Let $D_{\text{sft}}(x,y)=\mu(x)\times p_{\text{sft}}(y|x)$ be the SFT data distribution. Then, the SFT task is
${\theta}^{*}_{\text{sft}}=\arg\min_{\theta}\mathcal{L}_{\text{sft}}(\theta)\quad\quad\text{where}\quad\quad\mathcal{L}_{\text{sft}}(\theta):=E_{(x,y)\sim D_{\text{sft}}}\{A(\theta;x)-\theta^{\top}g(x,y)\},$ (19)
@@
$E_{x\sim\mu}\{\nabla_{\theta}A(\theta^{*}_{\text{sft}})\}=E_{(x,y)\sim D_{\text{sft}}}g(x,y).$ (20)
@@
Let $r$ be a reward function that determines the reward for each $(x,y).$ Let $\mathcal{L}_{\text{ro}}(\theta):=E_{x\sim\mu}E_{y\sim\pi_{\theta}}r(x,y)$. Then,
@@
where $D_{\text{KL}}(\pi_{\theta}\|{p})=E_{x\sim\mu}D_{\text{KL}}(\pi_{\theta}(\cdot|x)\|{p}(\cdot|x))$.
@@
$=E_{x\sim\mu}\{A(\theta;x)-A(\theta^{*}_{\text{sft}};x)-(\theta-\theta^{*}_{\text{sft}})^{\top}\nabla_{\theta}A(\theta^{*}_{\text{sft}};x)\}+\frac{1}{\beta}\mathcal{L}_{\text{ro}}(\theta)$ (24)
$=E_{x\sim\mu}\{A(\theta;x)-A(\theta^{*}_{\text{sft}};x)\}-(\theta-\theta^{*}_{\text{sft}})^{\top}E_{(x,y)\sim D_{\text{sft}}}g(x,y)+\frac{1}{\beta}\mathcal{L}_{\text{ro}}(\theta)$ (25)
@@
Thus, effectively this proves that the KL-RL objective enables multi-tasking between the SFT stage and the reward optimization RL objective. One may wonder why we did not pose the KL divergence regularizer on the transformed distributions through $D_{\text{KL}}(T(\pi)(\cdot\mid{\bm{x}})\|T({\pi_{\rm ref}})(\cdot|{\bm{x}}))$ instead. Consider the Best-of-$N$ jailbreaking for example. While the adversary may be using the model to generate $N$ responses and choose the least safe one for jailbreaking, the model should possess the core capabilities for other types of inference-time usage for other tasks that is different from that of jailbreaking (e.g., through chain-of-thought). Therefore, changing the KL divergence regularizer does not capture the fact that the model should remain suitable for all other tasks, and not just for the one for which it is getting aligned. We also note that if we used $D_{\text{KL}}(T(\pi)(\cdot\mid{\bm{x}})\|T({\pi_{\rm ref}})(\cdot|{\bm{x}}))$ instead, the problem would actually simplify to the standard alignment problem through a simple change of variables.
Appendix C Zero-shot prompt
We also evaluate the win rate using a prompted PaLM-XL model with the following instructions:
You are a helpful assistant, that ranks AI assistants’ responses by the quality of their answers. The AI assistants try to be helpful, sophisticated, emotionally aware, and humble-but-knowledgeable. Below are a series of dialogues between various people and an AI assistant, and the assistant tries to reply to the dialogue. I want you to rank the responses of assistants. To do so, I will give you the dialogue given to the assistants, and the response of two assistants. Please rank the assistants based on which response would be more helpful, sophisticated, emotionally aware, and humble-but-knowledgeable. All inputs are python dictionaries. Here is the prompt: { ”dialogue”: {dialogue}, } Here are the outputs of the assistants: [ { ”assistant”: ”assistant_1”, ”answer”: {output_1} }, { ”assistant”: ”assistant_2”, ”answer”: {output_2} } ] Respond 1 or 2 to indicate the better output. Please provide the ranking that the majority of humans would give.
Appendix D The effect of anchor points on calibration
In this section, we analyze the effect of using more anchor points on results. First, we present the full algorithm for more than one anchor point in Algorithm 3.
Algorithm 3 Offline Approximate Calibration with arbitrary anchor points
1:Base policy ${\pi_{\rm ref}}$, (uncalibrated) reward model $r$, training prompts ${\bm{x}}\in\mathcal{X}$, number of quantiles $p$, number of samples per prompt $m$.
2:for all ${\bm{x}}\in\mathcal{X}$ do
3:     Roll out $m$ i.i.d. samples from ${\pi_{\rm ref}}$ per training prompt ${\bm{x}}$: ${\bm{z}}_{j},\,j\in\{1...m\}$.
4:     for all $\ell\in\{0...p-1\}$ do
5:         $q_{\ell}=\frac{\ell+1}{p+1}$
6:         $r_{q_{\ell}}({\bm{x}})\leftarrow\text{top}_{q_{\ell}\cdot m}(r({\bm{x}},{\bm{z}}_{1}),\ldots r({\bm{x}},{\bm{z}}_{m}))$select $(q_{\ell}\cdot m)^{\text{th}}$ smallest element
7:     end for
8:end for
9:Define the calibrated reward function $\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}}):\mathcal{X}\times\mathcal{Y}\to[0,1]$ as follows:
${\log\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})}=\log\bigl{(}r({\bm{x}},{\bm{y}})\bigr{)}\cdot\frac{{\log\bigl{(}\frac{q_{a}+q_{b}}{2}}\bigr{)}}{\log\bigl{(}\frac{r_{q_{a}}({\bm{x}})+r_{q_{b}}({\bm{x}})}{2}\bigr{)}},$
where $a$ is the smallest number in $[0,p-1)$ such that $r_{q_{a}}({\bm{x}})\geq r({\bm{x}},{\bm{y}})$, otherwise $a=0$. Similarly, $b$ is the largest number in $[0,p-1)$ such that $r_{q_{b}}({\bm{x}})\leq r({\bm{x}},{\bm{y}})$, otherwise $b=p-1$.
10:Output: Calibrated Reward Function: $\widehat{\mathcal{C}}_{r,{\pi_{\rm ref}}}$ that assigns a calibrated reward for all prompts in the training set.
D.1 Better calibration leads to better gains
The effectiveness of calibration can be better understood when we approximate the calibration with fewer number of rollouts: $n$. When $n=1$, we take one other rollout from the base policy model and use that as a reference reward, similar to Wang et al. (2024), and with $n=3,5,7$, we approximate the median with this limited sample, and do the reward calibration. In Fig 5(a), we find that calibration using more number of samples improves the win rate, with $n=1$ providing no gains in the KL vs win rate tradeoff. Additionally, in Fig 5(b), when we calibrate with more anchor points, we find that the KL vs win rate tradeoffs improves as compared to the median-based calibration. This is further supported by the fact that with more anchor points, the calibration error is significantly reduced as shown in Fig 2(b). Hence, we find that as the calibration error is reduced, we achieve better KL vs win rate tradeoffs, thus able to better implicitly approximate the best-of-n model.
Figure 5: (a) Calibration improves KL vs win rate tradeoff even when we approximate calibration of median approximated with n=7 samples, whereas with n=1,3,5, the gains in win rate for a fixed KL-budget is lower, but still outperforms the uncalibrated model at higher KL values. Using n=100 samples is (calibrate_median_pareto) provides the best tradeoff (b) Calibration with two or more anchor points per-prompt improves KL-winrate tradeoff, as compared to median-based approximation - however there is diminishing gains with ten anchor points as compared to using two anchor points. The win rate is measured using the PaLM-2 M reward model trained on the Anthropic helpfulness preference dataset.
D.2 Calibration Reduces Reward Hacking

### v5

@@ B.2 Proof of Theorem 1
Proof of Eq. 2 The proof follows from the following identities:
$W^{\mathcal{T}}_{r}(\pi_{1}\succ\pi_{2}\mid{\bm{x}})$ $=\mathbb{E}_{{\bm{z}}\sim\mathcal{T}_{\pi_{1}}(\cdot\mid{\bm{x}}),{\bm{y}}\sim\mathcal{T}_{\pi_{2}}(\cdot\mid{\bm{x}})}\left\{w_{r}({\bm{z}},{\bm{y}}\mid{\bm{x}})\right\}$
$=\mathbb{E}_{{\bm{z}}\sim\mathcal{T}_{\pi_{1}}(\cdot\mid{\bm{x}})}\left\{\mathbb{E}_{{\bm{y}}\sim\mathcal{T}_{\pi_{2}}(\cdot\mid{\bm{x}})}\left\{w_{r}({\bm{z}},{\bm{y}}\mid{\bm{x}})\right\}\right\}$
$=\mathbb{E}_{{\bm{z}}\sim\mathcal{T}_{\pi_{1}}(\cdot\mid{\bm{x}})}\left\{\mathcal{C}_{r,\mathcal{T}_{\pi_{2}}}({\bm{x}},{\bm{z}})\right\}$
$=\sum_{{\bm{z}}}\mathcal{C}_{r,\mathcal{T}_{\pi_{2}}}({\bm{x}},{\bm{z}})\mathcal{T}_{\pi_{1}}({\bm{z}}\mid{\bm{x}}).$
@@
$\mathcal{L}(\pi(\cdot\mid{\bm{x}}),\alpha)=W^{\mathcal{T}}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})-\beta D_{\text{KL}}(\pi(\cdot\mid{\bm{x}})\|{\pi_{\rm ref}}(\cdot|{\bm{x}}))+\alpha\left(\sum_{{\bm{y}}}\pi({\bm{y}}\mid{\bm{x}})-1\right).$
@@
$\mathbf{0}=\frac{\partial\mathcal{L}(\pi(\cdot\mid{\bm{x}}),\alpha)}{\partial\pi({\bm{y}}\mid{\bm{x}})}=\frac{\partial W^{\mathcal{T}}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})}{\partial\pi({\bm{y}}\mid{\bm{x}})}-\beta\left(\log\frac{\pi({\bm{y}}\mid{\bm{x}})}{{\pi_{\rm ref}}({\bm{y}}\mid{\bm{x}})}+1\right)+\alpha.$
@@
$\mathcal{R}({\bm{x}},{\bm{y}})=\frac{\partial W^{\mathcal{T}}_{r}(\pi\succ{\pi_{\rm ref}}\mid{\bm{x}})}{\partial\pi({\bm{y}}\mid{\bm{x}})},$
@@
It remains to prove Eq. 7. Plugging in $\pi_{1}=\pi$, $\pi_{2}={\pi_{\rm ref}}$ to Eq. 2, and taking partial derivative with respect to $\pi({\bm{y}}\mid{\bm{x}})$ on the right hand side completes the proof. ∎
B.3 Proof of Lemma 2
@@ B.5 Proof of Lemma 4
Recall the definition of ${\bm{y}}_{u,{\bm{x}}}$ in Eq. 13, we have that
$\text{Pr}_{{\bm{y}}\sim{\pi_{\rm ref}}}\left(\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\leq u\right)$ $=\text{Pr}_{{\bm{y}}\sim{\pi_{\rm ref}}}\left(r({\bm{x}},{\bm{y}})\leq r({\bm{x}},{\bm{y}}_{u,{\bm{x}}})\right)$
@@
B.6 Proof of Lemma 5
For continuous language models, we have $\texttt{BoN}_{\pi}$ satisfies that $\forall{\bm{x}},{\bm{y}}$
$\text{Pr}_{{\bm{z}}\sim\texttt{BoN}_{\pi}(\cdot\mid{\bm{x}})}\left(r({\bm{x}},{\bm{y}})\leq r({\bm{x}},{\bm{y}})\right)=\text{Pr}_{{\bm{z}}\sim\pi(\cdot\mid{\bm{x}})}\left(r({\bm{x}},{\bm{z}})\leq r({\bm{x}},{\bm{y}})\right)^{N}=\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})^{N}.$ (15)
Hence
$\texttt{BoN}_{\pi}({\bm{y}}\mid{\bm{x}})=\frac{{\rm d}\text{Pr}_{{\bm{z}}\sim\texttt{BoN}_{\pi}(\cdot\mid{\bm{x}})}\left(r({\bm{x}},{\bm{z}})\leq r({\bm{x}},{\bm{y}})\right)}{{\rm d}{\bm{y}}}=\pi({\bm{y}}\mid{\bm{x}})\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})^{N-1}.$
For $\texttt{WoN}_{\pi}$, we have
$\text{Pr}_{{\bm{z}}\sim\texttt{WoN}_{\pi}(\cdot\mid{\bm{x}})}\left(r({\bm{x}},{\bm{y}})\leq r({\bm{x}},{\bm{y}})\right)=1-\text{Pr}_{{\bm{z}}\sim\pi(\cdot\mid{\bm{x}})}\left(r({\bm{x}},{\bm{z}})>r({\bm{x}},{\bm{y}})\right)^{N}=1-\left(1-\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\right)^{N}.$ (16)
Hence
$\texttt{WoN}_{\pi}({\bm{y}}\mid{\bm{x}})=\frac{{\rm d}\text{Pr}_{{\bm{z}}\sim\texttt{WoN}_{\pi}(\cdot\mid{\bm{x}})}\left(r({\bm{x}},{\bm{z}})\leq r({\bm{x}},{\bm{y}})\right)}{{\rm d}{\bm{y}}}=\pi({\bm{y}}\mid{\bm{x}})\left(1-\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\right)^{N-1}.$
B.7 Proof of Theorem 2
@@
If $\mathcal{T}$ is a calibrated inference-time procedure with mapping function $g_{\mathcal{T}}$, for any continuous language model $\pi$, $\beta>0$ and reward transformation function $\Phi$, we have that
$W^{\mathcal{T}}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})=\frac{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{\mathcal{T}}(F_{\Phi,\beta}(u))\int_{0}^{u}g_{\mathcal{T}}(u^{\prime}){\rm d}u^{\prime}{\rm d}u}{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{\mathcal{T}}(F_{\Phi,\beta}(u)){\rm d}u\int_{0}^{1}g_{\mathcal{T}}(u){\rm d}u}$
@@
$\mathcal{C}_{r,{\pi_{\rm ref}}}\circ\pi^{*}_{\mathcal{R}_{\Phi},\beta}(u\mid{\bm{x}})=\frac{e^{\Phi(u)/\beta}}{\int_{0}^{1}e^{\Phi(u^{\prime})/\beta}{\rm d}u^{\prime}}.$
@@
After the inference-time procedure $\mathcal{T}$ is applied, we have that inference-time base policy satisfies
$\mathcal{C}_{r,{\pi_{\rm ref}}}\circ\mathcal{T}_{{\pi_{\rm ref}}}(u\mid{\bm{x}})=\frac{g_{\mathcal{T}}(u)}{\int_{0}^{1}g_{\mathcal{T}}(u^{\prime}){\rm d}u^{\prime}}.$
@@
$\mathcal{C}_{r,{\pi_{\rm ref}}}\circ\mathcal{T}_{\pi^{*}_{\mathcal{R}_{\Phi},\beta}}(u\mid{\bm{x}})=\frac{\exp\left(\Phi(u)/\beta\right)g_{\mathcal{T}}(F_{\Phi,\beta}(u))}{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{\mathcal{T}}(F_{\Phi,\beta}(u)){\rm d}u}.$
Thus, the inference-time win rate satisfies
$W^{\mathcal{T}}_{r}(\pi^{*}_{\mathcal{R}_{\Phi},\beta}\succ{\pi_{\rm ref}}\mid{\bm{x}})$ $=\mathbb{E}_{{\bm{y}}\sim\mathcal{T}_{\pi^{*}_{\mathcal{R}_{\Phi},\beta}}(\cdot\mid{\bm{x}}),{\bm{z}}\sim\mathcal{T}_{{\pi_{\rm ref}}}(\cdot\mid{\bm{x}})}\left\{\mathbbm{1}\!\left\{r({\bm{x}},{\bm{y}})\geq r({\bm{x}},{\bm{z}})\right\}\right\}$
$=\mathbb{E}_{{\bm{y}}\sim\mathcal{T}_{\pi^{*}_{\mathcal{R}_{\Phi},\beta}}(\cdot\mid{\bm{x}}),{\bm{z}}\sim\mathcal{T}_{{\pi_{\rm ref}}}(\cdot\mid{\bm{x}})}\left\{\mathbbm{1}\!\left\{\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{y}})\geq\mathcal{C}_{r,{\pi_{\rm ref}}}({\bm{x}},{\bm{z}})\right\}\right\}$
$=\text{Pr}_{u\sim\mathcal{C}_{r,{\pi_{\rm ref}}}\circ\mathcal{T}_{\pi^{*}_{\mathcal{R}_{\Phi},\beta}}(\cdot\mid{\bm{x}}),u^{\prime}\sim\mathcal{C}_{r,{\pi_{\rm ref}}}\circ\mathcal{T}_{{\pi_{\rm ref}}}(\cdot\mid{\bm{x}})}\left(u^{\prime}\leq u\right)$
$=\frac{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{\mathcal{T}}(F_{\Phi,\beta}(u))\int_{0}^{u}g_{\mathcal{T}}(u^{\prime}){\rm d}u^{\prime}{\rm d}u}{\int_{0}^{1}\exp\left(\Phi(u)/\beta\right)g_{\mathcal{T}}(F_{\Phi,\beta}(u)){\rm d}u\int_{0}^{1}g_{\mathcal{T}}(u){\rm d}u},$
where the second equality follows from Lemma 2, completing the proof. ∎
B.8 Proof of Theorem 3
@@ B.9 Proof of Corollary 2
Let ${\bm{\mathcal{Y}}}=[0,1]$. We assume the LMs and reward models are context-independent. We use $u\in[0,1]$ to denote $\bm{y}$} and set the reward model to be $r(u)=u$. The base policy is a simple uniform distribution over $[0,1]$, ${\pi_{\rm ref}}=\text{Unif}([0,1])$. Note that $F_{\pi}(u)$ be the CDF of $\pi$, then we have that the BoN win rate is
@@
$=-N^{2}\int_{0}^{1}v^{N-1}F_{\pi}(v)^{N-1}\frac{\partial F_{\pi}(v)}{\partial\pi(u)}{\rm d}v$
@@
Appendix C The role of KL divergence in model alignment
@@
$\pi_{\theta}({\bm{y}}|{\bm{x}})=e^{\theta^{T}g({\bm{x}},{\bm{y}})-A(\theta;{\bm{x}})},$ (17)
where $g({\bm{x}},{\bm{y}})$ is a fixed encoding of $({\bm{x}},{\bm{y}})$, and $A(\theta;{\bm{x}})$ is the partition function normalizing the distribution.
@@
Let $D_{\text{sft}}({\bm{x}},{\bm{y}})=\mu({\bm{x}})\times p_{\text{sft}}(y|{\bm{x}})$ be the SFT data distribution. Then, the SFT task is
${\theta}^{*}_{\text{sft}}=\arg\min_{\theta}\mathcal{L}_{\text{sft}}(\theta)\quad\quad\text{where}\quad\quad\mathcal{L}_{\text{sft}}(\theta):=E_{({\bm{x}},{\bm{y}})\sim D_{\text{sft}}}\{A(\theta;{\bm{x}})-\theta^{\top}g({\bm{x}},{\bm{y}})\},$ (18)
@@
$E_{x\sim\mu}\{\nabla_{\theta}A(\theta^{*}_{\text{sft}})\}=E_{({\bm{x}},{\bm{y}})\sim D_{\text{sft}}}g({\bm{x}},{\bm{y}}).$ (19)
@@
Let $r$ be a reward function that determines the reward for each $({\bm{x}},{\bm{y}}).$ Let $\mathcal{L}_{\text{ro}}(\theta):=-E_{{\bm{x}}\sim\mu}E_{{\bm{y}}\sim\pi_{\theta}}r({\bm{x}},{\bm{y}})$. Then,
@@
where $D_{\text{KL}}(\pi_{\theta}\|{p})=E_{{\bm{x}}\sim\mu}D_{\text{KL}}(\pi_{\theta}(\cdot|{\bm{x}})\|{p}(\cdot|{\bm{x}}))$.
@@
$=E_{x\sim\mu}\{A(\theta;{\bm{x}})-A(\theta^{*}_{\text{sft}};{\bm{x}})-(\theta-\theta^{*}_{\text{sft}})^{\top}\nabla_{\theta}A(\theta^{*}_{\text{sft}};{\bm{x}})\}+\frac{1}{\beta}\mathcal{L}_{\text{ro}}(\theta)$ (23)
$=E_{x\sim\mu}\{A(\theta;{\bm{x}})-A(\theta^{*}_{\text{sft}};{\bm{x}})\}-(\theta-\theta^{*}_{\text{sft}})^{\top}E_{({\bm{x}},{\bm{y}})\sim D_{\text{sft}}}g({\bm{x}},{\bm{y}})+\frac{1}{\beta}\mathcal{L}_{\text{ro}}(\theta)$ (24)
@@
Thus, effectively this proves that the RLHF objective enables multi-tasking between the SFT stage and the reward optimization RL objective.
@@
One may wonder why we did not pose the KL divergence regularizer on the transformed distributions through $D_{\text{KL}}(\mathcal{T}_{\pi}(\cdot\mid{\bm{x}})\|\mathcal{T}_{{\pi_{\rm ref}}}(\cdot|{\bm{x}}))$ instead. Consider the Best-of-$N$ jailbreaking for example. While the adversary may be using the model to generate $N$ responses and choose the least safe one for jailbreaking, the model should possess the core capabilities for other types of inference-time usage for other tasks that is different from that of jailbreaking (e.g., through chain-of-thought). Therefore, changing the KL divergence regularizer does not capture the fact that the model should remain suitable for all other tasks, and not just for the one for which it is getting aligned. We also note that if we used $D_{\text{KL}}(\mathcal{T}_{\pi})(\cdot\mid{\bm{x}})\|\mathcal{T}_{{\pi_{\rm ref}}}(\cdot|{\bm{x}}))$ instead, the problem would actually simplify to the standard alignment problem through a simple change of variables.
Appendix D Calibration Reduces Reward Hacking